# The Assouad dimension of Kakeya sets in $\mathbb{R}^{3}$

### v1

@@ 4.1 Arrangements of rectangular prisms
Recall that $\tilde{S}$ is a rectangular prism of dimensions roughly $t\theta\times t\times t$. Define $r=t\theta$; then by pigeonholing, we can find a ball $B\subset\tilde{S}$ of radius $r$ so that
$\Big{|}B\cap\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\gtrapprox\delta^{3\eta}\ |B|.$
By $\delta_{0}$ sufficiently small depending on $\eta$ and the implicit constants in the above quasi-inequality, we can ensure that
$\Big{|}B\cap\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\geq\delta^{4\eta}|B|,$ (4.12)
and since $\theta\geq\delta^{-4\eta/\varepsilon}\frac{s}{t}$, we have $\delta^{4\eta}\geq(s/r)^{\varepsilon}$. Thus (4.12) implies (4.4), and hence Conclusion (A) holds.
Step 4. Suppose instead that for every prism $R_{0}\in\mathcal{R}$ and every $\theta\geq\delta^{-4\eta/\varepsilon}\frac{s}{t}$, (4.6) fails. Let $s_{1}=\delta^{-4\eta/\varepsilon}s$. Since $s\leq\delta^{\varepsilon}t$, if we select $\eta$ sufficiently small then $s_{1}\leq t$. Select a $\gtrapprox\delta^{\eta}$ dense shading $Y_{3}(R)\subset Y_{2}(R)$ so that $\operatorname{P-spread}(x)\leq\frac{s_{1}}{t}$ for each $x\in Y_{3}(R)$. Let $\mathcal{R}_{0}\subset\mathcal{R}$ consist of those $R\in\mathcal{R}$ with $|Y_{3}(R)|\gtrapprox\delta^{\eta}|R|$; we will choose the implicit constant and implicit power of $|\log\delta|$ so that #$\#\mathcal{R}_{0}\gtrapprox\delta^{\eta}(\#\mathcal{R}$).
For each $x\in\bigcup_{R\in\mathcal{R}_{0}}Y_{3}(R)$, define the “line spread”
$\operatorname{L-spread}(x)=\max\Big{\{}\angle\big{(}\ell(R),\ell(R^{\prime})\big{)}\Big{\}},$ (4.13)
where the maximum is taken over all pairs $R,R^{\prime}$ for which $x\in Y_{3}(R)$ and $x\in Y_{3}(R^{\prime})$, and where $\angle(\ell,\ell^{\prime})$ denotes the unsigned angle between the lines $\ell$ and $\ell^{\prime}$. Since $\ell(R)$ and $\ell(R^{\prime})$ are defined up to uncertainty $O(t)$, $\operatorname{L-spread}(x)$ is meaningfully defined if it is substantially larger than $t$. If $x\not\in\bigcup_{R\in\mathcal{R}_{0}}Y_{3}(R)$, define $\operatorname{L-spread}(x)=0$.
We claim there is a number $\theta\in[\delta^{-\varepsilon^{2}/2}t,\pi/4]$ so that the shading
$Y_{4}(R)=\{x\in Y_{3}(R)\colon\operatorname{L-spread}(x)\in[\theta,2\theta)\}$ (4.14)
is $\gtrapprox\delta^{\eta}$ dense. We argue as follows: if this was not the case, then
$\sum_{R\in\mathcal{R}_{0}}|\{x\in Y_{3}(R)\colon$ $\operatorname{L-spread}(x)<\delta^{-\varepsilon^{2}/2}t\}|\gtrapprox\sum_{R\in\mathcal{R}_{0}}|Y_{3}(R)|$
$\gtrapprox\delta^{2\eta}\sum_{R\in\mathcal{R}}|Y_{1}(R)|\gtrsim\delta^{2\eta}\mu\Big{|}\bigcup_{R\in\mathcal{R}}Y_{1}(R)\Big{|}$
$\geq\delta^{2\eta}\mu\Big{|}\bigcup_{R\in\mathcal{R}_{0}}Y_{3}(R)\Big{|},$
and hence there must exist a point $x\in\bigcup_{R\in\mathcal{R}_{0}}Y_{3}(R)$ with
#$\#\{R\in\mathcal{R}_{0}\colon x\in Y_{3}(R)\}\gtrapprox\delta^{2\eta}\mu\geq\delta^{-\varepsilon^{2}+5\eta}(st\#\mathcal{R}),\ \ \operatorname{L-spread}(x)\leq\delta^{-\varepsilon^{2}/2}t,\ \ \operatorname{P-spread}(x)\leq\frac{s_{1}}{t}.$ (4.15)
@@
(the last condition follows from the fact that $x\in\bigcup Y_{3}(R)$). But this means that the rectangular prisms $\{R\in\mathcal{R}_{0}\colon x\in Y_{3}(R)\}$ are contained in the $O(1)$ dilate of a rectangular prism of dimensions $s_{1}\times\delta^{-\varepsilon^{2}/2}t\times 1$. Since the prisms in $\mathcal{R}$ satisfy the Convex Wolff Axioms with error $\delta^{-\eta}$, we have
#$\#\{R\in\mathcal{R}_{0}\colon x\in Y_{3}(R)\}\lesssim\delta^{-\eta}\Big{(}s_{1}\cdot\delta^{-\varepsilon^{2}/2}t\Big{)}(\#\mathcal{R})=\delta^{-3\eta/\varepsilon-\eta-\varepsilon^{2}/2}(st\#\mathcal{R}),$ (4.16)
Comparing (4.15) and (4.16), we obtain a contradiction if $\delta_{0}$ and $\eta$ are sufficiently small.
Step 5. At this point we can fix $\theta\geq\delta^{-\varepsilon^{2}/2}t$, and we may suppose that the corresponding shading $Y_{4}$ from (4.14) is $\gtrapprox\delta^{\eta}$ dense. For each prism $R\in\mathcal{R}_{0}$, let $\hat{R}$ be the prism of dimensions $\frac{s_{1}\theta}{t}\times\theta\times 1$ obtained by anisotropically dilating $R$. We claim that if $R_{0},R\in\mathcal{R}_{0}$ are prisms and if $x\in Y_{4}(R_{0})\cap Y_{3}(R)$, then $R\subset 2\hat{R}_{0}$. This follows from the fact that $\angle(\Pi(R_{0}),\Pi(R))\leq\frac{s_{1}}{t}$, and $\angle(\ell(R_{0}),\ell(R))\leq 2\theta$. In fact more is true—if we define $w=\max\big{(}\frac{s_{1}\theta}{t},t\big{)}$, then $N_{w}(R)$ is comparable to a prism of dimensions $\frac{s_{1}\theta}{t}\times w\times 1$, and this prism is contained in $2\hat{R}_{0}$. Furthermore, $N_{w}(R)\cap R_{0}$ is comparable to a rectangular prism of dimensions $s\times t\times w/\theta$.
@@
Let $\mathcal{R}_{1}\subset\mathcal{R}_{0}$ consist of those rectangular prisms for which $|Y_{4}(R)|\gtrapprox\delta^{2\eta}|R|$. Let $R_{0}\in\mathcal{R}_{1}$; recall that $R_{0}$ is a $s\times t\times 1$ rectangular prism; for notational simplicity we will suppose $R_{0}=[-s/2,s/2]\times[-t/2,t/2]\times[-1/2,1/2]$, and that the intersection of $Y_{4}(R_{0})$ with the $z$-axis has measure $\gtrapprox\delta^{\eta}$. Define $\tilde{R}_{0}=[-\frac{s_{1}\theta}{t},\frac{s_{1}\theta}{t}]\times[-\theta,\theta]\times[-1,1]$.
@@
Select $\frac{w}{\theta}$-separated points $z_{1},\ldots,z_{N}\in Y_{4}(R_{0})$ on the $z$-axis, with $N\gtrapprox\delta^{\eta}\frac{\theta}{w}$. We will label the points so that $|z_{i}-z_{j}|\geq\frac{w}{\theta}|i-j|$. For each index $i$, select a prism $R_{i}\in\mathcal{R}_{0}$ with $z_{i}\in Y_{3}(R_{i})$ and $\angle((0,0,1),\ell(R))\in[\theta,2\theta)$.
@@
For each index $i$, let $R_{i}^{\prime}$ be the $\frac{s_{1}\theta}{t}$ neighborhood of $R_{i}$, and let $Y_{3}^{\prime}(R_{i}^{\prime})=N_{\frac{s_{1}\theta}{t}}(Y_{3}(R_{i}))$. Since $R_{i}\in\mathcal{R}_{0}$ (and hence $|Y_{3}(R_{i})\gtrapprox\delta^{\eta}|R_{i}|$), we have
$|Y_{3}^{\prime}(R_{i}^{\prime})|\gtrapprox\delta^{\eta}|R_{i}^{\prime}|.$ (4.17)
@@
$R_{i}^{\prime}$ is comparable to a prism of dimensions $\frac{s_{1}\theta}{t}\times w\times 1$ whose coaxial line makes angle $\sim\theta$ with the $z$ axis. We have
$|R_{i}^{\prime}\cap R_{j}^{\prime}|\lesssim\big{(}\frac{s_{1}\theta}{t}\big{)}\frac{w}{1+|i-j|},$
@@
$\Big{\|}\sum_{i=1}^{N}\chi_{Y_{3}^{\prime}(R_{i}^{\prime})}\Big{\|}_{2}^{2}\leq\sum_{i,j=1}^{N}|R_{i}^{\prime}\cap R_{j}^{\prime}|\lesssim N\frac{s_{1}\theta w}{t}\big{|}\log(\frac{\theta}{w}\big{)}\big{|}.$
@@
$\Big{|}\bigcup_{i=1}^{N}Y_{3}^{\prime}(R_{i}^{\prime})\Big{|}\geq\frac{\Big{(}\sum|Y_{3}^{\prime}(R_{i}^{\prime})|\Big{)}^{2}}{\Big{\|}\sum\chi_{Y_{3}^{\prime}(R_{i}^{\prime})}\Big{\|}_{2}^{2}}\gtrapprox\frac{\Big{(}N\delta^{\eta}\frac{s_{1}\theta}{t}w\Big{)}^{2}}{N\frac{s_{1}\theta w}{t}}=N\delta^{2\eta}\frac{s_{1}\theta}{t}w\gtrapprox\delta^{3\eta}\frac{s_{1}}{t}.$ (4.18)
On the other hand, each set $R_{i}^{\prime}$ is contained in $\hat{R}_{0}$, and thus
$\Big{|}\hat{R}_{0}\cap N_{\frac{s_{1}\theta}{t}}\Big{(}\bigcup_{\begin{subarray}{c}R\in\mathcal{R}\\ R\subset\hat{R}_{0}\end{subarray}}Y(R)\Big{)}\Big{|}\gtrapprox\delta^{3\eta}|\hat{R}_{0}|.$ (4.19)
Recall that $\hat{R}_{0}$ is a prism of dimensions $2\frac{s_{1}\theta}{t}\times 2\theta\times 2$, and $s_{1}=\delta^{-3\eta/\varepsilon}s$. Define $s^{\prime}=\frac{s_{1}\theta}{t}$ and $t^{\prime}=\min(1,t\frac{s^{\prime}}{s})=\min(1,\delta^{-3\eta/\varepsilon}\theta)$. Then $s^{\prime}\in[\delta^{-\varepsilon^{2}/2}s,\frac{s_{1}}{t}]$; since $s\leq\delta^{\varepsilon}t$, we can select $\eta$ sufficiently small so that $\frac{s_{1}}{t}\leq\delta^{\varepsilon/2}$, and hence $s^{\prime}\in[\delta^{-\varepsilon^{2}/2}s,\delta^{\varepsilon/2}]$.
By (4.19) we conclude that for each $R_{0}\in\mathcal{R}_{1}$, if $\hat{R}_{0}^{\prime}$ is the corresponding $s^{\prime}\times t^{\prime}\times 1$ prism and $\eta$ is selected sufficiently small compared to $\varepsilon$ and $\tau$, then the shading $\hat{Y}^{\prime}(\hat{R}_{0}^{\prime})$ given by (4.5) satisfies
$|\hat{Y}^{\prime}(\hat{R}_{0}^{\prime})|\gtrapprox\delta^{3\eta}|\hat{R}_{0}|\geq\delta^{3\eta+3\eta/\varepsilon}|\hat{R}_{0}^{\prime}|\geq\delta^{\tau}|\hat{R}_{0}^{\prime}|.$ (4.20)
@@
=$\mathcal{R}^{\prime}=\{\hat{R}_{0}^{\prime}\colon\hat{R}_{0}\in\mathcal{R}_{1}\}.$
(4.20) ensures that $|Y^{\prime}(R^{\prime})|\geq\delta^{\tau}|R^{\prime}|$ for each $R^{\prime}\in\mathcal{R}^{\prime}.$ Since $\mathcal{R}$ satisfies the Convex Wolff axioms with error $\delta^{-\eta}$ and #$\#\mathcal{R}_{1}\gtrapprox\delta^{2\eta}\#\mathcal{R}_{1}$, we have that $\mathcal{R}_{1}$ satisfies the Convex Wolff axioms with error $\lessapprox\delta^{-3\eta}$, and hence by Remark 2.7 we have that $\mathcal{R}^{\prime}$ satisfies the Convex Wolff Axioms with error $\lessapprox\delta^{-3\eta}$; if $\delta_{0}$ and $\eta$ is selected sufficiently small then this is $\leq\delta^{-\tau}$. ∎
The culmination of Section 4.1 is the following result, which is obtained by iterating Lemma 4.4.
@@
Then $\bigcup_{R\in\mathcal{R}}Y(R)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\eta}$.
@@
Fix $\varepsilon>0$ and let let $\eta,\delta_{0}$ be small quantities to be determined below. Let $\delta\in(0,\delta_{0}]$, $s,t\in[\delta,1]$ with $s\leq\delta^{\varepsilon}t$; let $\mathcal{R}$ be a set of $s\times t\times 1$ prisms that satisfy the Convex Wolff Axioms with error $\delta^{-\eta}$; and let $Y(R)$ be a $\delta^{\eta}$-dense shading.
Define $\mathcal{R}_{1}=\mathcal{R}$ and define the shading $Y_{1}=Y$. Let $s_{1}=s$ and $t_{1}=t$. Let $\eta=\tau_{1}<\tau_{2}<\tau_{3}<\ldots<\tau_{N},$ $N=\lfloor 2/\varepsilon^{2}\rfloor$ be small quantities to be specified below. We will perform the following iterative process for each $i=1,2,\ldots,N$. At the beginning of step $i$, the following things will be true
@@
Apply Lemma 4.4 to $(\mathcal{R}_{i},Y_{i})$ with $\varepsilon$ as above and $\tau_{i+1}$ in place of $\tau$. If $\tau_{i}>0$ is sufficiently small depending on $\varepsilon$ and $\tau_{i+1}$, then the Lemma can be applied (indeed, $\tau_{i}$ must be less than or equal to the quantity “$\eta$” from the conclusion of Lemma 4.4). The quantity $\delta_{0}$ (and hence $\delta$) will be chosen sufficiently small so that Lemma 4.4 can be applied at each step.

### v2

@@ 4.1 Arrangements of rectangular prisms
Recall that $S^{\dagger}$ is a rectangular prism of dimensions roughly $\tilde{t}\theta\times\tilde{t}\times\tilde{t}$.
Next, undo the unit rescaling with respect to $2T_{u}$; recall that this unit rescaling anisotropically dilated two directions by a multiplicative factor of roughly $u^{-1}$, and left the third direction unchanged. Thus the preimage of $S^{\dagger}$ under this rescaling is comparable to a rectangular prisms whose smallest dimension is at least $\gtrsim\tilde{t}\theta u=t\theta\geq\delta^{-7\varepsilon_{1}/\varepsilon}s$.
By pigeonholing, we can find a ball $B$ of radius $r\gtrsim\delta^{-7\varepsilon_{1}/\varepsilon}s$ contained in the preimage of $S^{\dagger}$, so that
$\Big{|}B\cap\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\gtrapprox\delta^{6\varepsilon_{1}}|B|\geq\delta^{-\varepsilon_{1}}(s/r)^{\varepsilon}|B|.$ (4.16)
If $\delta_{0}$ is selected sufficiently small, then the LHS of (4.16) is at least $(s/r)^{\varepsilon}|B|$, and thus Conclusion (A) holds.
Step 4: Ensuring that $u$ is large. At this point we may assume that for every $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$, every prism $\tilde{R}_{0}\in\tilde{\mathcal{R}}^{T_{u}}$, and every $\theta\geq\delta^{-7\varepsilon_{1}/\varepsilon}(\tilde{s}/\tilde{t})$, we have that (4.10) fails.
For each $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$ and each $\tilde{R}\in\tilde{\mathcal{R}}^{T_{u}}$, define
$\tilde{Y}_{5}(\tilde{R})=\{x\in\tilde{Y}_{4}(R)\colon\operatorname{P-spread}(x)\leq\delta^{-7\varepsilon_{1}/\varepsilon}(\tilde{s}/\tilde{t}).$
For each $R\in\bigcup_{T_{u}\in\mathbb{T}_{u}^{\prime\prime}}\mathcal{R}[2T_{u}]$, define $Y_{5}(R)\subset Y_{4}(R)$ to be the preimage of $\tilde{Y}_{5}(\tilde{R})$ under the unit rescaling associated to $2T_{u}$. Define $Y_{5}(R)=\emptyset$ if $R\not\in\bigcup_{T_{u}\in\mathbb{T}_{u}^{\prime\prime}}\mathcal{R}[2T_{u}]$. We have that
  • For each $T_{u}\in\mathbb{T}_{u},$
    $\sum_{R\in\mathcal{R}[2T_{u}]}|Y_{5}(R)|\gtrsim\sum_{R\in\mathcal{R}[2T_{u}]}|Y_{4}(R)|\gtrsim\delta^{2\varepsilon_{1}}\sum_{R\in\mathcal{R}[2T_{u}]}|R|.$ (4.17)
  • $\sum_{R\in\mathcal{R}}|Y_{5}(R)|\gtrsim\sum_{R\in\mathcal{R}}|Y_{4}(R)|\gtrsim\delta^{2\varepsilon_{1}}\sum_{R\in\mathcal{R}}|R|.$
  • For each $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$ and each $x\in\bigcup_{R\in\mathcal{R}[2T_{u}]}Y_{5}(R)$, we have
    $\mu\leq\#\{R\in\mathcal{R}[2T_{u}]\colon x\in Y_{5}(R)\}<2\mu.$ (4.18)
We claim that
$u\geq\delta^{-\varepsilon^{2}/4}t.$ (4.19)
To verify this claim, fix a choice of $T_{u}\in\mathbb{T}_{u}^{\prime\prime}$ and a choice of $x\in\bigcup_{R\in\mathcal{R}[2T_{u}]}Y_{5}(R)$. The prisms in the set (4.18) are contained in a prism of volume $\lesssim\delta^{-7\varepsilon_{1}/\varepsilon}su^{2}/t$; this is because the unit rescaling of these prisms are contained in a rectangular prism of dimensions comparable to $\delta^{-7\varepsilon_{1}/\varepsilon}(\tilde{s}/\tilde{t})\times 1\times 1$. This latter set has volume roughly $\delta^{-7\varepsilon_{1}/\varepsilon}s/t$, and undoing the unit rescaling distorts volume by a factor of $u^{2}$.
@@
On the other hand by (4.19) and (4.4), at least $\mu\geq\delta^{-(2/3)\varepsilon^{2}}(st\#\mathcal{R})$ prisms from $\mathcal{R}$ must be contained in this set. Since $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $\delta^{-\eta}$, we conclude that
$\delta^{-(2/3)\varepsilon^{2}}(st\#\mathcal{R})\leq\delta^{-\eta}(\delta^{-7\varepsilon_{1}/\varepsilon}su^{2}/t)(\#\mathcal{R}).$
re-arranging, we obtain (4.19) provided we select $\eta$ and $\varepsilon_{1}$ sufficiently small.
Step 5: Constructing larger prisms. Define $s_{1}=\delta^{-7\varepsilon_{1}/\varepsilon}s$. Since $s\leq\delta^{\varepsilon}t$, if we select $\varepsilon_{1}$ sufficiently small (depending on $\varepsilon$), then $s_{1}\leq t$. For each prism $R\in\mathcal{R}_{0}$, let $\hat{R}$ be the prism of dimensions $\frac{s_{1}u}{t}\times u\times 1$ obtained by anisotropically dilating $R$. We claim that if $R_{0},R\in\mathcal{R}$ are prisms and if $x\in Y_{5}(R_{0})\cap Y_{4}(R)$, then $R\subset 2\hat{R}_{0}$. We verify this claim as follows: we have that $R_{0}$ and $R$ are contained in the 2-fold dilate of a common $u$-tube $T_{u}$, and furthermore $\angle(n(\tilde{R}_{0}),n(\tilde{R}))\leq s_{1}/t$ (here $\tilde{R}_{0}$ and $\tilde{R}$ are the unit rescalings of $R_{0}$ and $R$ with respect to $2T_{u}$). This means that $\tilde{R}$ is contained in the $s_{1}/t$ neighbourhood of the plane $\Pi(\tilde{R}_{0})$; but after undoing the unit rescaling with respect to $2T_{u}$, the preimage of $N_{s_{1}/t}(B(0,1)\cap\Pi(\tilde{R}_{0}))$ is comparable to $\hat{R}$.
In fact a superficially stronger statement is true — if we define $w=\max\big{(}\frac{s_{1}u}{t},t\big{)}$, then $N_{s_{1}u/t}(R)$ is comparable to a prism of dimensions $\frac{s_{1}u}{t}\times w\times 1$, and this prism is contained in $2\hat{R}_{0}$. Furthermore, $N_{s_{1}u/t}(R)\cap R_{0}$ is comparable to a rectangular prism of dimensions $s\times t\times w/u$.
@@
Let $\mathcal{R}_{0}\subset\mathcal{R}$ consist of those rectangular prisms for which $|Y_{5}(R)|\geq c_{1}\delta^{2\varepsilon_{1}}|R|$. We will choose $c_{1}\sim 1$ sufficiently small so that $\sum_{R\in\mathcal{R}_{0}}|Y_{5}(R)|\gtrsim\sum_{R\in\mathcal{R}}|Y_{5}(R)|$. Let $R_{0}\in\mathcal{R}_{0}$. Recall that $R_{0}$ is a $s\times t\times 1$ rectangular prism; for notational simplicity we will suppose $R_{0}=[-s/2,s/2]\times[-t/2,t/2]\times[-1/2,1/2]$, and that the intersection of $Y_{5}(R_{0})$ with the $z$-axis has measure $\gtrapprox\delta^{\eta}$. Then $2\hat{R}_{0}=[-\frac{s_{1}u}{t},\frac{s_{1}u}{t}]\times[-u,u]\times[-1,1]$.
@@
Select $\frac{w}{u}$-separated points $z_{1},\ldots,z_{N}\in Y_{5}(R_{0})$ on the $z$-axis, with $N\gtrsim\delta^{2\varepsilon_{1}}\frac{u}{w}$. We will label the points so that $|z_{i}-z_{j}|\geq\frac{w}{u}|i-j|$. For each index $i$, select a prism $R_{i}\in\mathcal{R}$ with $z_{i}\in Y_{4}(R_{i})$ and $\angle((0,0,1),\ell(R))\in[c_{2}u,u)$. If we select $c_{2}\sim 1$ sufficiently small, then such a prism $R_{i}$ always exists by (4.8) and (4.9).
@@
For each index $i$, let $R_{i}^{\dagger}$ be the $\frac{s_{1}u}{t}$ neighborhood of $R_{i}$, and let $Y_{1}^{\dagger}(R_{i}^{\dagger})=N_{\frac{s_{1}u}{t}}(Y_{1}(R_{i}))$. Since $|Y_{1}(R_{i})|\geq\frac{1}{2}\delta^{\eta}|R_{i}|$, we have
$|Y_{1}^{\dagger}(R_{i}^{\dagger})|\gtrsim\delta^{\eta}|R_{i}^{\dagger}|.$ (4.20)
@@
$R_{i}^{\dagger}$ is comparable to a prism of dimensions $\frac{s_{1}u}{t}\times w\times 1$ whose coaxial line makes angle $\sim u$ with the $z$ axis. We have
$|R_{i}^{\dagger}\cap R_{j}^{\dagger}|\lesssim\big{(}\frac{s_{1}u}{t}\big{)}\frac{w}{1+|i-j|},$
@@
$\Big{\|}\sum_{i=1}^{N}\chi_{Y_{1}^{\dagger}(R_{i}^{\dagger})}\Big{\|}_{2}^{2}\leq\sum_{i,j=1}^{N}|R_{i}^{\dagger}\cap R_{j}^{\dagger}|\lesssim N\frac{s_{1}uw}{t}|\log(u/w)|.$
@@
$\Big{|}\bigcup_{i=1}^{N}Y_{1}^{\dagger}(R_{i}^{\dagger})\Big{|}\geq\frac{\Big{(}\sum|Y_{1}^{\dagger}(R_{i}^{\dagger})|\Big{)}^{2}}{\Big{\|}\sum\chi_{Y_{1}^{\dagger}(R_{i}^{\dagger})}\Big{\|}_{2}^{2}}\gtrapprox\frac{\Big{(}N\delta^{2\varepsilon_{1}}\frac{s_{1}u}{t}w\Big{)}^{2}}{N\frac{s_{1}uw}{t}}=N\delta^{4\varepsilon_{1}}\frac{s_{1}u}{t}w\gtrapprox\delta^{6\varepsilon_{1}}\frac{s_{1}}{t}u^{2}.$ (4.21)
On the other hand, each set $R_{i}^{\prime}$ is contained in $\hat{R}_{0}$ (a set of volume roughly $s_{1}u^{2}/t$), and thus
$\Big{|}2\hat{R}_{0}\cap N_{\frac{s_{1}u}{t}}\Big{(}\bigcup_{\begin{subarray}{c}R\in\mathcal{R}\\ R\subset\hat{R}_{0}\end{subarray}}Y(R)\Big{)}\Big{|}\gtrapprox\delta^{6\varepsilon_{1}}|2\hat{R}_{0}|.$ (4.22)
Recall that $2\hat{R}_{0}$ is a prism of dimensions $2\frac{s_{1}u}{t}\times 2u\times 2$, and $s_{1}=\delta^{-7\varepsilon_{1}/\varepsilon}s$. Define $s^{\prime}=\frac{s_{1}u}{t}$ and $t^{\prime}=\min(1,t\frac{s^{\prime}}{s})=\min(1,\delta^{-7\varepsilon_{1}/\varepsilon}u)$. By (4.19), we have $s^{\prime}\in[\delta^{-\varepsilon^{2}/4}s,\frac{s_{1}}{t}]$. Since $s\leq\delta^{\varepsilon}t$, we can select $\eta$ sufficiently small so that $\frac{s_{1}}{t}\leq\delta^{\varepsilon/2}$, and hence $s^{\prime}\in[\delta^{-\varepsilon^{2}/4}s,\delta^{\varepsilon/2}]$.
By (4.22) we conclude that for each $R_{0}\in\mathcal{R}_{0}$, if $\hat{R}_{0}^{\prime}$ is the corresponding $s^{\prime}\times t^{\prime}\times 1$ prism and $\varepsilon_{1}$ is selected sufficiently small compared to $\varepsilon$ and $\tau$, then the shading $\hat{Y}^{\prime}(\hat{R}_{0}^{\prime})$ given by (4.2) satisfies
$|\hat{Y}^{\prime}(\hat{R}_{0}^{\prime})|\gtrapprox\delta^{6\varepsilon_{1}}|\hat{R}_{0}|\geq\delta^{6\varepsilon_{1}+7\varepsilon_{1}/\varepsilon}|\hat{R}_{0}^{\prime}|\geq\delta^{\tau}|\hat{R}_{0}^{\prime}|.$ (4.23)
@@
=$\mathcal{R}^{\prime}=\{\hat{R}_{0}^{\prime}\colon\hat{R}_{0}\in\mathcal{R}_{0}\}.$
(4.23) ensures that $|Y^{\prime}(R^{\prime})|\geq\delta^{\tau}|R^{\prime}|$ for each $R^{\prime}\in\mathcal{R}^{\prime}.$ Since $\mathcal{R}$ satisfies the Convex Wolff axioms with error $\delta^{-\eta}$ and #$\#\mathcal{R}_{0}\gtrapprox\delta^{2\varepsilon_{1}}\#\mathcal{R}$, we have that $\mathcal{R}_{0}$ satisfies the Convex Wolff axioms with error $\lessapprox\delta^{-2\varepsilon_{1}-\eta}$, and hence by Remark 2.7 we have that $\mathcal{R}^{\prime}$ satisfies the Convex Wolff Axioms with error $\lessapprox\delta^{-2\varepsilon_{1}-\eta}$. If $\varepsilon_{1},\eta,$ and $\delta_{0}$ and $\eta$ is selected sufficiently small, then this is $\leq\delta^{-\tau}$. ∎
Soon, we will iterate Lemma 4.2 multiple times. If Conclusion (B) happens repeatedly, then eventually our planks will have width 1, i.e. our planks will become rectangular slabs of dimensions $s\times 1\times 1$ for some $s<\!\!<1$. The next lemma says that the union of such a collection of slabs has volume close to 1.
Lemma 4.4.
Let $20<s\leq 1/$ 2 and $K\geq 1$. Let $\mathcal{R}$ be a multi-set of $s\times 1\times 1$ rectangular prisms, and suppose that $\mathcal{R}$ satisfies the Convex Wolff Axioms with error $K$. Let $\{Y(R),R\in\mathcal{R}\}$ be a $K^{-1}$-dense shading. Then
$\Big{|}\bigcup_{R\in\mathcal{R}}Y(R)\Big{|}\gtrsim K^{-3}(\log 1/s)^{-1}.$ (4.24)
Proof.
$K^{-1}s(\#\mathcal{R})\leq\int\chi_{E}\cdot\sum_{R\in\mathcal{R}}\chi_{Y(R)}\leq\int\chi_{E}\cdot\sum_{R\in\mathcal{R}}\chi_{R}\leq|E|^{1/2}\Big{\|}\sum_{R\in\mathcal{R}}\chi_{R}\Big{\|}_{2}.$ (4.25)
In the argument that follows, we will obtain an upper bound for $\big{\|}\sum_{\mathcal{R}}\chi_{R}\big{\|}_{2}$.
For each prism $R\in\mathcal{R}$, there is an associated normal vector $v(R)\in S^{2}$, which is well-defined up to uncertainty $O(s)$. For $R,R^{\prime}\in\mathcal{R}$, define $\angle(R,R^{\prime})$ to be the (unsigned) angle between $v(R)$ and $v(R^{\prime})$.
Fix $R\in\mathcal{R}$. Observe that if $R\cap R^{\prime}$ is non-empty and $\angle(R,R^{\prime})=\theta\sim 2^{j}s$, then $R^{\prime}$ is contained in the $O(\theta)\times 1\times 1$ rectangular prism concentric with $R$. Since $\mathcal{R}$ obeys the Convex Wolff Axioms, at most $K(2^{j}s)(\#\mathcal{R})$ slabs $R^{\prime}\in\mathcal{R}$ can have this property.
Let $j_{1}\sim\log(1/s)$ be the smallest integer with $2^{j_{1}}s\geq 1$. Using the above observation, we compute
$\begin{split}\Big{\|}\sum_{R\in\mathcal{R}}\chi_{R}\Big{\|}_{2}^{2}&=\sum_{R^{\prime}\in\mathcal{R}}|R\cap R^{\prime}|=\sum_{j=0}^{j_{1}}\sum_{\begin{subarray}{c}R^{\prime}\colon R\cap R^{\prime}\neq\emptyset,\\ \angle(R,R^{\prime})\sim 2^{j}s\end{subarray}}|R\cap R^{\prime}|\\ &\lesssim\sum_{j=0}^{j_{1}}\Big{(}K(2^{j}s\#\mathcal{R})\Big{)}\Big{(}\frac{s^{2}}{2^{j}s}\Big{)}\lesssim Ks^{2}(\#\mathcal{R})\log(1/s).\end{split}$ (4.26)
Chaining (4.25) and (4.26), we obtain (4.24). ∎
The culmination of Section 4.1 is the following result, which says that every arrangement of planks satisfying the Convex Wolff Axioms must have discretized Assouad dimension close to 3.
@@
Then $\bigcup_{\mathcal{R}}Y(R)$ has discretized Assouad dimension at least $3-\varepsilon$ at scale $\delta$ and scale separation $\delta^{-\eta}$.
In brief, we prove Proposition 4.5 by iterating Lemma 4.2. If Conclusion (A) holds at any step in the iteration, then our union of planks has discretized Assouad dimension close to 3, and we are done. If Conclusion (B) holds, then we repeat the process with the resulting wider collection of planks. Eventually, our planks will have width 1, at which point we can apply Lemma 4.4 to again conclude that our union of planks has discretized Assouad dimension close to 3. We now turn to the details.
@@
Fix $\varepsilon>0$ and let $\eta,\delta_{0}$ be small quantities to be determined below. Let $\delta\in(0,\delta_{0}]$, $s,t\in[\delta,1]$ with $s\leq\delta^{\varepsilon}t$; let $\mathcal{R}$ be a set of $s\times t\times 1$ prisms that satisfy the Convex Wolff Axioms with error $\delta^{-\eta}$; and let $Y(R)$ be a $\delta^{\eta}$-dense shading.
Define $\mathcal{R}_{1}=\mathcal{R}$ and define the shading $Y_{1}=Y$. Let $s_{1}=s$ and $t_{1}=t$. Let $\eta=\tau_{1}<\tau_{2}<\tau_{3}<\ldots<\tau_{N},$ $N=\lfloor 4/\varepsilon^{2}\rfloor$ be small quantities to be specified below. We will perform the following iterative process for each $i=1,2,\ldots,N$. At the beginning of step $i$, the following properties will hold:
  • (a)
@@
When $i=1$, the hypotheses of Proposition 4.5 guarantee that the above properties hold.
Suppose that the above properties hold at step $i$. Apply Lemma 4.2 to $(\mathcal{R}_{i},Y_{i})$ with $\varepsilon$ as above and $\tau_{i+1}$ in place of $\tau$. If $\tau_{i}>0$ is sufficiently small depending on $\varepsilon$ and $\tau_{i+1}$, then the Lemma can be applied (indeed, $\tau_{i}$ must be less than or equal to the quantity “$\eta$” from the conclusion of Lemma 4.2). The quantity $\delta_{0}$ (and hence $\delta$) will be chosen sufficiently small so that Lemma 4.2 can be applied at each step.